OpenStudy (anonymous):

Average:1332 Standard diviation:725 (a) What proportion of tax returns show a refund greater than $2,000? (b) What proportion of the tax returns show that the taxpayer owes money to the government? (c) What proportion of the tax returns show a refund between $100 and $700?

OpenStudy (amistre64):

lol ... i need the practice :)

OpenStudy (amistre64):

z = (2000 - 1332)/725 = .9214

OpenStudy (amistre64):

what can you find me for a z of .92?

OpenStudy (anonymous):

.8212?

OpenStudy (amistre64):

i agree. .8212 is everything less than our zscore; we want whats greater than it sooo: 1 - .8212 should be good

OpenStudy (amistre64):

the total is 1 and we want whats left over

OpenStudy (anonymous):

.1788 for a

OpenStudy (amistre64):

good

OpenStudy (amistre64):

so when the refund is less than 0 they owe right?

OpenStudy (amistre64):

z=(0 - 1332)/725 = ?

OpenStudy (anonymous):

-1.8372

OpenStudy (amistre64):

z = -1.84 is what i get as well

OpenStudy (amistre64):

so what value can you get me for a z of -1.84?

OpenStudy (anonymous):

.9671?

OpenStudy (amistre64):

thats a bit high; but the table is measureing zs that are to the right of the mean; so subtract .5 from it to begin with

OpenStudy (amistre64):

OR, 1 - .9671 should do it since that is what it would be if we flipped the chart

OpenStudy (anonymous):

ok so .0329

OpenStudy (amistre64):

.0329 is good

OpenStudy (amistre64):

the trick is in figuring out the ztables info and where we need to be in relation to it

OpenStudy (amistre64):

next one: $100 and $700 z1 = (100-1332)/725 z2 = (700-1332)/725

OpenStudy (amistre64):

both of these are to the left of the mean so they are -z values

OpenStudy (anonymous):

-1.6993 and -.8717

OpenStudy (amistre64):

those look good, and the table is gonna give us the ... "wrong" values for both of them so we need to take the table and subtract it FROM 1

OpenStudy (amistre64):

then we can determine the distance between z2 and z1 by subtracting the "corrected" values

OpenStudy (anonymous):

.0455 and .1922?

OpenStudy (amistre64):

.0475 and .1922 is what i get

OpenStudy (anonymous):

so it would be .1467?

OpenStudy (amistre64):

.1447 if you adjust for your typo ;)

OpenStudy (amistre64):

or is that my typo?

OpenStudy (amistre64):

-1.70 the round off right?

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=z%3D-1.699

OpenStudy (anonymous):

oh it is .1476

OpenStudy (amistre64):

.0447

OpenStudy (anonymous):

I didn't round to -1.70 earlier

OpenStudy (amistre64):

.1475 is in the ball park

OpenStudy (anonymous):

thank you for helping me

OpenStudy (amistre64):

yep, thanks for the practice :)